# Semi-supervised Underwater Image Enhancement Using A Physics-Aware Triple-Stream Network

### v3

@@ V-A Underwater Image Enhancement
We evaluated the proposed PATS-UIENet along with 16 baselines using two full-reference quantitative metrics and two non-reference quantitative metrics. A qualitative analysis was also performed. In this subsection, the results obtained in the three experiments will be reported.
@@ V-A1 Full-Reference Quantitative Evaluation
Since the Test-U80, Test-S110, and Test-Scenes testing sets contain reference images, a full-reference quantitative evaluation was conducted on the proposed method and the 16 baselines using the PSNR and SSIM metrics across these testing sets. The results are shown in Table I. As can be seen, our PATS-UIENet achieved the best result in terms of both the metrics no matter which testing set was used. Compared with the prior-based baselines using IFM [1], such as DCP [3], UDCP [4], GDCP [5] and DHL [8], our method showed a significant improvement.
| Method | Test-U80 | Test-S110 | Test-Scenes | |||
| PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |
| DCP | 12.48 | 63.34 | 14.52 | 67.38 | 18.09 | 68.95 |
| UDCP | 9.37 | 33.62 | 10.04 | 34.28 | 14.76 | 56.43 |
| GDCP | 12.46 | 67.41 | 14.67 | 71.29 | 14.11 | 63.50 |
| DHL | 14.49 | 62.81 | 14.96 | 69.68 | 15.16 | 53.92 |
| HE | 16.48 | 77.77 | 15.37 | 74.76 | 13.61 | 62.82 |
| Retinex | 10.38 | 61.25 | 8.29 | 53.44 | 9.50 | 55.56 |
| MMLE | 18.56 | 76.21 | 17.32 | 77.39 | 14.89 | 62.23 |
| HLRP | 13.56 | 22.76 | 12.55 | 29.24 | 12.12 | 18.21 |
| WaterNet | 17.02 | 70.27 | 18.80 | 75.19 | 25.84 | 84.02 |
| UColor | 20.93 | 85.19 | 20.23 | 84.23 | 24.31 | 81.40 |
| FUnIE-GAN | 15.31 | 59.68 | 16.89 | 66.32 | 25.11 | 83.16 |
| USUIR | 16.92 | 68.69 | 17.54 | 76.58 | 15.14 | 66.51 |
| PUIE-Net (MP) | 22.26 | 88.89 | 21.85 | 89.19 | 21.85 | 73.47 |
| PUIE-Net (MC) | 22.05 | 88.80 | 21.72 | 89.20 | 20.85 | 73.10 |
| U-Transformer | 20.98 | 72.76 | 19.92 | 68.25 | 24.09 | 79.98 |
| URanker | 21.87 | 85.75 | 21.27 | 87.43 | 22.46 | 82.35 |
| PATS-UIENet (Ours) | 23.01 | 89.89 | 22.48 | 90.26 | 27.20 | 88.30 |
@@ V-A2 Non-reference Quantitative Evaluation
We used two non-reference metrics, i.e., UIQM [55] and UCIQE [56], to evaluate the performance of our method and sixteen baselines on the Test-U80, Test-S110, Test-Scenes, Test-C60, Test-UCCS and Test-R53 testing sets. The results are shown in Tables II. It can be seen that the performance of our method was lower than the best counterpart but was comparable to most baselines. As demonstrated in [22, 27, 35, 26], the UIQM and UCIQE metrics are not always consistent with the quality of images processed by UIE methods. These findings may explain the relatively lower UIQM and UCIQE values produced by our method.
@@
| UIQM | UCIQE | UIQM | UCIQE | UIQM | UCIQE | UIQM | UCIQE | UIQM | UCIQE | UIQM | UCIQE | |
| DCP | 1.35 | 0.51 | 1.36 | 0.57 | 1.61 | 0.57 | 0.56 | 0.51 | 0.03 | 0.44 | -2.05 | 0.42 |
| UDCP | 2.69 | 0.59 | 1.23 | 0.60 | 2.80 | 0.60 | 1.24 | 0.54 | 1.83 | 0.57 | 0.76 | 0.59 |
| GDCP | 2.17 | 0.56 | 2.12 | 0.64 | 1.89 | 0.65 | 1.49 | 0.60 | 1.24 | 0.56 | 5.50 | 0.50 |
| DHL | 3.19 | 0.64 | 2.35 | 0.66 | 3.54 | 0.66 | 1.85 | 0.64 | 2.98 | 0.63 | 0.22 | 0.64 |
| HE | 3.58 | 0.67 | 2.97 | 0.68 | 4.27 | 0.67 | 2.32 | 0.65 | 3.93 | 0.67 | 1.71 | 0.64 |
| Retinex | 2.33 | 0.48 | 0.85 | 0.45 | 3.36 | 0.47 | 1.21 | 0.46 | 2.03 | 0.47 | -2.03 | 0.51 |
| MMLE | 3.89 | 0.58 | 2.72 | 0.60 | 3.60 | 0.61 | 2.77 | 0.57 | 4.06 | 0.57 | 3.18 | 0.55 |
| HLRP | 3.71 | 0.63 | 2.88 | 0.65 | 3.85 | 0.66 | 1.98 | 0.65 | 3.89 | 0.64 | 1.98 | 0.65 |
| WaterNet | 2.11 | 0.50 | 1.47 | 0.54 | 3.67 | 0.57 | 1.28 | 0.51 | 1.25 | 0.45 | 0.09 | 0.45 |
| UColor | 3.44 | 0.60 | 2.30 | 0.63 | 3.64 | 0.57 | 2.08 | 0.58 | 3.22 | 0.55 | 1.28 | 0.54 |
| FUnIE-GAN | 3.73 | 0.50 | 3.59 | 0.57 | 4.43 | 0.58 | 4.50 | 0.51 | 3.47 | 0.49 | 0.19 | 0.42 |
| USUIR | 4.10 | 0.63 | 2.84 | 0.65 | 3.88 | 0.65 | 3.12 | 0.62 | 4.19 | 0.60 | 3.31 | 0.60 |
| PUIE-Net (MP) | 3.42 | 0.56 | 2.25 | 0.59 | 3.54 | 0.57 | 1.82 | 0.55 | 4.03 | 0.52 | 1.56 | 0.51 |
| PUIE-Net (MC) | 3.31 | 0.56 | 2.22 | 0.59 | 3.11 | 0.57 | 1.78 | 0.55 | 3.82 | 0.52 | 1.30 | 0.49 |
| U-Transformer | 4.37 | 0.55 | 3.68 | 0.56 | 4.54 | 0.55 | 4.19 | 0.53 | 4.19 | 0.53 | 2.43 | 0.53 |
| URanker | 3.64 | 0.57 | 2.55 | 0.59 | 3.68 | 0.57 | 2.36 | 0.56 | 4.17 | 0.56 | 2.25 | 0.56 |
| PATS-UIENet (Ours) | 3.76 | 0.57 | 2.47 | 0.60 | 3.70 | 0.57 | 2.24 | 0.56 | 3.58 | 0.55 | 2.55 | 0.55 |
@@ V-A3 Qualitative Analysis
The visual results generated by the sixteen baselines and our method on six testing sets are shown in Figs. 4 - 9, respectively. It can be seen that some methods with the higher UIQM [55] or UCIQE [56] score did not produce visually pleased results. For example, MMLE [14] increased the contrast but produced images with pale colors (see Fig. 4). FUnIE-GAN [20] could not completely remove the color cast in degraded images (see Fig. 7). HE[9], HLRP[15], and USUIR[24] introduced artifacts that reduced the visual quality (see Fig. 8). In contrast, our method not only effectively removed the color cast from degraded images but also significantly improved the visibility of the images. It is demonstrated that the images that our method enhanced still exhibit the satisfactory visual quality even if the UIQM or UCIQE value that our method generated is lower than that produced by the baselines. This observation is consistent with the findings presented in the literature [22, 27, 35, 26].
@@ V-B Transmission Estimation and Color Restoration
The quantitative evaluation of our methods and four prior-based baselines for transmission estimation and color restoration is reported in Table III. It can be seen that our method achieved the superior, or at least the comparable, performance to that of the four prior-based baselines on four different subsets of Test-R53. Specifically, our method outperformed the baselines with a large margin on both the Michmoret and Katzaa subsets, no matter which metric was considered. It should be noted that none of these methods produced a PCC value higher than 0.12 on the Satil subset. The inferior results may be attributed to the special scenes contained in this subset [35]. As shown in Fig. 10, our method was able to perform transmission estimation and color restoration by directly learning from the limited real-world underwater images.
@@
| DCP | -0.16 | 34.63 | 0.00 | 35.47 | -0.26 | 34.75 | 0.03 | 36.21 |
| UDCP | -0.54 | 36.97 | -0.15 | 40.43 | 0.06 | 38.80 | 0.12 | 51.45 |
| GDCP | 0.29 | 33.76 | 0.24 | 34.47 | -0.04 | 34.28 | -0.03 | 35.31 |
| DHL | -0.07 | 32.10 | 0.10 | 35.50 | 0.36 | 35.92 | 0.11 | 34.82 |
| Ours | 0.61 | 11.68 | 0.32 | 9.47 | -0.38 | 16.87 | 0.02 | 10.36 |
@@ V-C1 Effect of the Semi-supervised Learning Framework
To further validate the effect of semi-supervised learning framework, we provide a quantitative comparison among the supervised learning, bi-directional supervised learning, and semi-supervised learning frameworks. As reported in Table IV, either the supervised learning framework or the bi-directional supervised learning framework normally generated the lower metric values than those produced by the proposed semi-supervised learning framework, in terms of the PSNR, SSIM, PCC and metrics. It has been suggested that our network could be better trained using the semi-supervised learning framework with both the labeled and unlabeled real-world images, which improved the enhancement performance.
| Learning Framework | Test-U80 | Michmoret | ||
| PSNR | SSIM | PCC | ||
| Sup | 21.09 | 86.72 | 0.51 | 17.86 |
| Bi-sup | 22.06 | 88.95 | 0.65 | 15.59 |
| Semi-sup (Ours) | 23.01 | 89.89 | 0.61 | 11.68 |
V-C2 Effect of the RCT and the RCM
To examine the effect of different modules on the performance of our PATS-UIENet, we removed the RCT, the RCM or both of them. In this case, we derived three variants of the PATS-UIENet: one with both RCT and RCM removed, one with only RCT removed and one with only RCM removed, denoted as “Simplified”, “w/o RCT” and “w/o RCM”, respectively. The results produced by these variants and the PATS-UIENet are reported in Table V. It can be observed that the PATS-UIENet with the entire components normally achieved the best performance in terms of the PSNR, SSIM, PCC and metrics. Although the utilization of the RCT had a negative effect on the PSNR value (observed by comparing “Simplified” with “w/o RCM”), the performance of transmission estimation was increased. It was indicated that the adaptability of the red channel increased the robustness and generalization of the model trained.
| Variant | Test-U80 | Michmoret | ||
| PSNR | SSIM | PCC | ||
| Simplified | 22.18 | 88.56 | 0.59 | 13.35 |
| w/o RCT | 22.71 | 89.43 | 0.60 | 11.42 |
| w/o RCM | 21.84 | 88.98 | 0.61 | 14.05 |
| PATS-UIENet | 23.01 | 89.89 | 0.61 | 11.68 |
V-C3 Effect of Different Network Depth
To evaluate the impact of the network depth on the performance of our PATS-UIENet, we changed the number of layers of the network from 4 to 3 and 5, individually. In total, we derived two variants of the PATS-UIENet with regard to different network depths. In terms of both the variants, the numbers of filters at each block were set to 64, 128, 256, 256, 128, 64 and 64 and 64, 128, 256, 512, 1024, 1024, 512, 256, 128, 64 and 64, respectively. As shown in Table VI, the proposed PATS-UIENet which comprised four layers produced the more promising results compared with the two variants.
| Network Depth | Test-U80 | Michmoret | ||
| PSNR | SSIM | PCC | ||
| 3 Layers | 22.43 | 89.40 | 0.61 | 13.29 |
| 5 Layers | 22.57 | 89.45 | 0.62 | 9.41 |
| 4 Layers (Ours) | 23.01 | 89.89 | 0.61 | 11.68 |
V-C4 Effect of the Revised IFM
To further investigate the effect of the revised IFM [2] on UIE, we used the outputs of the B-Stream and A-Stream of the proposed PATS-UIENet as the transmission map and the ambient light data, respectively, required by the original IFM [1], as shown in Eq. (1). In this case, the results obtained using the original IFM are compared with those produced by the revised IFM in Table VII. As can be seen, the revised IFM achieved the superior performance to the original IFM in terms of the PSNR, SSIM and metrics. It is indicated that the revised IFM models the underwater degradation process better than the orignal IFM.

### v4

@@ V-A Underwater Image Enhancement
We evaluated the proposed PATS-UIENet along with sixteen baselines using three full-reference quantitative metrics and two non-reference quantitative metrics. A qualitative analysis and a performance analysis were also performed. In this subsection, the results obtained in the four experiments will be reported.
@@ V-A1 Full-Reference Quantitative Evaluation
Since the Test-U80, Test-S110 and Test-Scenes testing sets contain reference images, a full-reference quantitative evaluation was conducted on the proposed method and the 16 baselines using the PSNR, SSIM [57] and LPIPS [58] metrics across these testing sets. The results are shown in Table I. As can be seen, our PATS-UIENet normally achieved the best performance across all three metrics, regardless of the testing set. Compared with the baselines designed on top of the IFM [1], such as UDCP [4] and DHL [8], and the existing semi-supervised learning approaches, including DDFormer [29], Semi-UIR [28] and UWFormer [30], our method normally showed a large margin.
| Method | Test-U80 | Test-S110 | Test-Scenes | ||||||
| PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |
| UDCP [4] | 9.51 | 33.66 | 41.74 | 10.07 | 34.29 | 37.66 | 14.76 | 56.43 | 30.31 |
| DHL [10] | 15.16 | 63.93 | 30.40 | 14.61 | 62.93 | 28.39 | 14.99 | 69.73 | 22.36 |
| HE [9] | 16.60 | 77.86 | 25.67 | 15.41 | 74.79 | 30.05 | 13.61 | 62.82 | 36.10 |
| HLRP [15] | 13.56 | 22.76 | 33.90 | 12.55 | 29.24 | 33.18 | 12.12 | 18.21 | 39.35 |
| MMLE [14] | 18.56 | 76.21 | 22.57 | 17.32 | 77.39 | 22.77 | 14.89 | 62.23 | 31.99 |
| WaterNet [22] | 17.13 | 70.41 | 40.01 | 18.83 | 75.22 | 31.08 | 25.84 | 84.02 | 16.28 |
| UColor [23] | 21.05 | 85.24 | 17.97 | 20.26 | 84.25 | 15.07 | 24.31 | 81.40 | 19.05 |
| FUnIE-GAN [20] | 15.31 | 59.68 | 44.28 | 16.89 | 66.32 | 34.98 | 25.11 | 83.16 | 13.76 |
| USUIR [24] | 16.92 | 68.69 | 28.63 | 17.54 | 76.58 | 21.38 | 15.14 | 66.51 | 32.44 |
| PUIE-Net (MP) [25] | 22.26 | 88.89 | 12.01 | 21.85 | 89.19 | 9.70 | 21.85 | 73.47 | 24.45 |
| U-Transformer [21] | 20.98 | 72.76 | 26.32 | 19.92 | 68.25 | 30.17 | 24.09 | 79.98 | 15.08 |
| URanker [26] | 21.87 | 85.75 | 18.64 | 21.27 | 87.43 | 13.51 | 22.46 | 82.35 | 20.87 |
| CCMSRNet [27] | 22.74 | 88.74 | 13.18 | 22.22 | 89.07 | 10.76 | 18.96 | 77.67 | 23.85 |
| DDFormer [29] | 10.94 | 25.74 | 58.26 | 10.87 | 41.27 | 72.19 | 10.64 | 27.81 | 69.46 |
| Semi-UIR [28] | 23.63 | 81.81 | 23.08 | 19.83 | 73.95 | 30.54 | 19.12 | 74.94 | 25.68 |
| UWFormer [30] | 19.65 | 85.26 | 17.21 | 21.45 | 90.27 | 10.88 | 19.62 | 80.36 | 21.05 |
| PATS-UIENet (Ours) | 23.59 | 90.16 | 10.42 | 22.78 | 90.54 | 8.85 | 25.87 | 84.34 | 12.04 |
@@ V-A2 Non-reference Quantitative Evaluation
We used two non-reference metrics, i.e., UIF [59] and MUSIQ [60], to assess the performance of our method and sixteen baselines on the Test-U80, Test-S110, Test-Scenes, Test-C60, Test-UCCS and Test-R53 testing sets. The results are shown in Table II. It can be seen that our method achieved the best UIF score on three out of the six testing sets, including Test-U80, Test-S110 and Test-C60, and ranked the second on the Test-R53 testing set. Regarding the MUSIQ metric, the proposed method still demonstrated the competitive performance, especially on the challenging Test-S110, Test-Scenes and Test-R53 testing sets, even though it did not always produce the highest score.
@@
| UIF | MUSIQ | UIF | MUSIQ | UIF | MUSIQ | UIF | MUSIQ | UIF | MUSIQ | UIF | MUSIQ | |
| UDCP [4] | 36.40 | 44.99 | 32.01 | 51.38 | 61.25 | 36.99 | 25.41 | 33.03 | 34.72 | 29.72 | 3.24 | 36.85 |
| DHL [10] | 41.32 | 47.43 | 14.80 | 57.22 | 0.55 | 36.57 | 25.04 | 37.27 | 3.62 | 31.07 | 2.53 | 42.64 |
| HE [9] | 44.04 | 46.68 | 45.01 | 56.14 | 52.81 | 35.79 | 41.50 | 36.08 | 38.08 | 31.24 | 1.67 | 43.06 |
| HLRP [15] | 0.59 | 49.73 | 0.91 | 56.87 | 0.71 | 40.21 | 3.57 | 34.65 | 0.32 | 34.44 | 1.26 | 43.53 |
| MMLE [14] | 30.99 | 52.50 | 35.05 | 60.18 | 42.05 | 44.67 | 24.95 | 40.18 | 28.08 | 35.69 | 24.67 | 53.53 |
| WaterNet [22] | 33.25 | 31.84 | 28.50 | 40.47 | 72.94 | 38.47 | 28.65 | 34.75 | 33.38 | 24.88 | 9.71 | 32.93 |
| UColor [23] | 44.15 | 45.10 | 42.83 | 54.80 | 70.20 | 35.12 | 26.46 | 36.01 | 49.22 | 35.12 | 2.31 | 49.63 |
| FUnIE-GAN [20] | 29.12 | 47.47 | 22.77 | 52.27 | 37.69 | 45.38 | 23.11 | 43.49 | 22.67 | 34.64 | 10.93 | 38.92 |
| USUIR [24] | 5.17 | 44.44 | 37.02 | 56.83 | 48.82 | 37.91 | 22.35 | 34.15 | 20.54 | 28.83 | 17.98 | 39.44 |
| PUIE-Net (MP) [25] | 9.18 | 49.05 | 60.98 | 60.44 | 27.69 | 45.46 | 3.39 | 38.69 | 51.76 | 29.18 | 46.16 | 45.15 |
| U-Transformer [21] | 5.16 | 39.34 | 24.94 | 44.96 | 38.68 | 30.44 | 29.31 | 37.66 | 29.69 | 27.05 | 19.24 | 30.06 |
| URanker [26] | 54.87 | 44.63 | 58.31 | 57.15 | 72.54 | 39.20 | 44.70 | 38.37 | 45.97 | 29.60 | 39.19 | 41.92 |
| CCMSRNet [27] | 54.61 | 49.85 | 0.23 | 60.59 | 63.59 | 39.64 | 41.85 | 40.24 | 51.34 | 33.72 | 34.80 | 48.23 |
| DDFormer [29] | -5.36 | 36.99 | 1.92 | 40.68 | -3.16 | 23.87 | 11.14 | 36.21 | -2.14 | 20.19 | 5.86 | 24.87 |
| Semi-UIR [28] | 50.06 | 45.40 | 31.94 | 51.22 | 44.01 | 41.47 | 46.82 | 43.25 | 32.28 | 30.34 | 23.33 | 35.21 |
| UWFormer [30] | 46.48 | 50.95 | 55.69 | 61.53 | 67.40 | 36.80 | 36.76 | 41.40 | 53.24 | 34.30 | 39.98 | 46.26 |
| PATS-UIENet (Ours) | 58.83 | 49.91 | 61.20 | 60.66 | 68.07 | 46.70 | 46.99 | 40.49 | 51.18 | 31.63 | 42.51 | 46.63 |
@@ V-A3 Qualitative Analysis
The enhanced images generated by the 16 baselines and our method on six testing sets are shown in Figs. 4 - 9, respectively. It can be seen that some methods with the higher UIF [59] or MUSIQ [60] score did not produce visually pleasing results. For example, MMLE [14] produced the highest MUSIQ score on Test-U80 and Test-R53, but the resulting images suffered from pale colors and over-enhancement artifacts (see Figs. 4 and 9). In contrast, our PATS-UIENet improved different degraded images and produced images with the natural and vivid color. Although the UIF [59] or MUSIQ [60] scores that our method produced were slightly lower than those obtained using some baseline methods [14, 22, 20] on certain testing sets, the enhanced images still manifest visually satisfactory quality.
V-A4 Performance Analysis
We compared the proposed PATS-UIENet with 11 deep baseline methods in terms of the number of parameters, Floating Point Operations Per Second (FLOPs) and inference speed (ms). Both the number of parameters and the FLOPs were calculated using the ptflops222https://pypi.org/project/ptflops/ tool. As shown in Table III, the values of number of parameters, FLOPs and inference speed vary significantly. Specifically, UColor [23] exhibits the largest number of parameters (148.04M) and the largest FLOPs value (2804.36G), while UWFormer [30] incurs the slowest inference speed (491.00 ms). On the other hand, USUIR [24] has the lightest design with only 0.23M parameters and has a FLOPs value of 29.62G and owns the fastest inference speed (2.13ms). In contrast, our PATS-UIENet achieved a proper balance between the number of parameters or FLOPs and the inference speed.
| Method | #Params (M) | FLOPs (G) | Speed (ms) |
| WaterNet [22] | 1.09 | 142.84 | 6.40 |
| UColor [23] | 148.04 | 2804.36 | 15.51 |
| FUnIE-GAN [20] | 7.02 | 20.48 | 73.18 |
| USUIR [24] | 0.23 | 29.62 | 2.13 |
| PUIE-Net (MP) [25] | 1.40 | 70.54 | 101.89 |
| U-Transformer [21] | 22.8 | 5.96 | 91.30 |
| URanker [26] | 3.15 | 20.90 | 45.50 |
| CCMSRNet [27] | 21.69 | 87.18 | 152.51 |
| DDFormer [29] | 7.63 | 35.54 | 50.04 |
| Semi-UIR [28] | 1.65 | 72.88 | 84.93 |
| UWFormer [30] | 29.84 | 30.18 | 491.00 |
| PATS-UIENet (Ours) | 45.71 | 179.70 | 86.17 |
@@ V-B Transmission Estimation and Color Restoration
The quantitative evaluation of our methods and four prior-based baselines for transmission estimation and color restoration is reported in Table IV. It can be seen that our method achieved the superior, or at least the comparable, performance to that of the four prior-based baselines on four different subsets of Test-R53. Specifically, our method outperformed the baselines with a large margin on both the Michmoret and Katzaa subsets, no matter which metric was considered. It should be noted that none of these methods produced a PCC value higher than 0.12 on the Satil subset. The inferior results may be attributed to the special scenes contained in this subset [39]. As shown in Fig. 10, our method was able to perform transmission estimation and color restoration by directly learning from the limited real-world underwater images.
@@
| DCP [3] | -0.16 | 34.63 | 0.00 | 35.47 | -0.26 | 34.75 | 0.03 | 36.21 |
| UDCP [4] | -0.54 | 36.97 | -0.15 | 40.43 | 0.06 | 38.80 | 0.12 | 51.45 |
| GDCP [5] | 0.29 | 33.76 | 0.24 | 34.47 | -0.04 | 34.28 | -0.03 | 35.31 |
| DHL [10] | -0.07 | 32.10 | 0.10 | 35.50 | 0.36 | 35.92 | 0.11 | 34.82 |
| Ours | 0.62 | 12.83 | 0.32 | 9.47 | -0.28 | 15.31 | 0.04 | 13.14 |
@@ V-C1 Effect of the Semi-supervised Learning Framework
To validate the effect of the proposed semi-supervised learning framework, we compared it with the unsupervised learning, supervised learning and bi-directional supervised learning frameworks. As reported in Table V, the proposed semi-supervised framework always outperformed the other frameworks in terms of different metrics across different testing sets. It has been suggested that our network can be better trained using the semi-supervised learning framework with both the labeled and unlabeled real-world images, which improves the enhancement performance.
| Learning Framework | Test-U80 | Michmoret | |||
| PSNR | SSIM | LPIPS | PCC | ||
| Unsupervised | 16.85 | 78.44 | 29.91 | 0.50 | 5.46 |
| Supervised | 22.19 | 88.23 | 13.07 | 0.57 | 17.01 |
| Bi-supervised | 23.00 | 89.49 | 11.93 | 0.61 | 15.02 |
| Semi-supervised (Ours) | 23.59 | 90.16 | 10.42 | 0.62 | 12.83 |
V-C2 Effect of the RCT and RCM
To examine the effect of the RCT and RCM on the performance of our PATS-UIENet, we removed the RCT, the RCM or both of them. In this case, we derived three variants of the PATS-UIENet: one with both the RCT and RCM removed, one with only the RCT removed and one with only the RCM removed, denoted as “Simplified”, “w/o RCT” and “w/o RCM”, respectively. As shown in Table VI, the complete PATS-UIENet achieved the best performance in terms of all the metrics. Removal of the RCT led to a slight increase in the PSNR and SSIM metrics compared to the Simplified version, but worsened the LPIPS and scores. On the other hand, removal of the RCM reduced the performance with regard to all metrics across the two data sets, confirming its usefulness in facilitating information exchange between the B-Stream and A-Stream. The results demonstrate that both the RCT and RCM contribute meaningfully to the performance of our method.
| Variant | Test-U80 | Michmoret | |||
| PSNR | SSIM | LPIPS | PCC | ||
| Simplified | 22.96 | 89.64 | 11.25 | 0.59 | 13.59 |
| w/o RCM | 22.10 | 88.83 | 12.20 | 0.60 | 13.50 |
| w/o RCT | 23.14 | 90.03 | 12.20 | 0.60 | 13.97 |
| Ours | 23.59 | 90.16 | 10.42 | 0.62 | 12.83 |
V-C3 Effect of the Degradation Control Factor
To evaluate the impact of the degradation control factor on the performance of the PATS-UIENet, we conducted an ablation experiment on testing three different values of . As shown in Table VII, the default setting of = 0.1 achieved the best performance with regard to both the full-reference and the non-reference metrics across the two data sets.
| Test-U80 | Michmoret | ||||
| PSNR | SSIM | LPIPS | PCC | ||
| 0.001 | 23.21 | 89.77 | 11.02 | 0.59 | 13.92 |
| 1 | 23.06 | 89.72 | 11.08 | 0.61 | 14.11 |
| 0.1 (Ours) | 23.59 | 90.16 | 10.42 | 0.62 | 12.83 |
V-C4 Effect of Loss Hyperparameters
To evaluate the impact of the hyperparameters used for the loss functions, including , , and , on the performance of the proposed PATS-UIENet, we conducted a comprehensive ablation study by changing their values. Regarding the supervised learning loss function (see Eq. (7)), the results produced by our method with different combinations of the and values are shown in Table VIII. It can be seen that the combination that we used produced the best result in terms of each metric. For the unsupervised learning loss function (see Eq. (15)), the results produced by our method with different values of are reported in Table IX. Again, the value that we chose led to the best result no matter what metric was considered. In terms of the semi-supervised learning loss function (see Eq. (16)), we present the results obtained using our method with different values in Table X. As can be observed, the value that we utilized produced the best result with regard to each metric. The above findings highlight the effectiveness of the values of different loss hyperparameters that we chose.
| Test-U80 | Michmoret | |||||
| PSNR | SSIM | LPIPS | PCC | |||
| 1 | 0.0005 | 22.04 | 88.01 | 12.94 | 0.36 | 16.92 |
| 0.0001 | 0.0005 | 22.84 | 89.65 | 10.85 | 0.61 | 13.57 |
| 0.1 | 0.1 | 22.95 | 89.67 | 11.52 | 0.61 | 15.69 |
| 0.1 | 0 | 22.91 | 80.57 | 11.54 | 0.59 | 16.92 |
| 0.1 | 0.0005 (Ours) | 23.59 | 90.16 | 10.42 | 0.62 | 12.83 |
| Test-U80 | Michmoret | ||||
| PSNR | SSIM | LPIPS | PCC | ||
| 0.1 | 22.95 | 89.80 | 11.01 | 0.61 | 15.14 |
| 10 | 22.36 | 88.19 | 13.09 | 0.60 | 14.72 |
| 1 (Ours) | 23.59 | 90.16 | 10.42 | 0.62 | 12.83 |
| Test-U80 | Michmoret | ||||
| PSNR | SSIM | LPIPS | PCC | ||
| 1 | 22.32 | 88.75 | 12.16 | 0.61 | 14.08 |
| 0.01 | 23.31 | 89.96 | 10.92 | 0.61 | 12.84 |
| 0.1 (Ours) | 23.59 | 90.16 | 10.42 | 0.62 | 12.83 |
V-C5 Effect of Different Stream Architectures
To further examine the effect of different stream architectures on the PATS-UIENet, we constructed its three variants by building the D-Stream, B-Stream and A-Stream using convolution and/or Transformer networks. In addition, we obtained a fourth variant by removing the D-Stream. In essence, this variant is equal to the network built on top of the original IFM [1]. The four variants were trained and tested using the same setup as that used for our PATS-UIENet. The results produced by these variants and our method are shown in Table XI. It can be seen that our method produced the better, or at least comparable, results in contrast to the four variants. This finding supports the design philosophy of the PATS-UIENet. That is to say, CNNs are suitable for capturing local characteristics while Transformer is good at encoding global characteristics. Moreover, our network which was built on top of the revised IFM [2] normally outperformed the variant constructed based on the original IFM [1], confirming the choice of the revised IFM [2].
| Method | Streams | Test-U80 | Michmoret | |||||
| D- | B- | A- | PSNR | SSIM | LPIPS | PCC | ||
| CNN | CNN | CNN | 22.31 | 88.69 | 12.48 | 0.58 | 12.44 | |
| Trans | Trans | CNN | 20.71 | 75.36 | 25.31 | 0.73 | 17.28 | |
| Trans | Trans | Trans | 20.57 | 75.00 | 25.73 | 0.71 | 14.23 | |
| N/A | CNN | Trans | 22.51 | 88.85 | 11.47 | 0.71 | 15.95 | |
| Ours | CNN | CNN | Trans | 23.59 | 90.16 | 10.42 | 0.62 | 12.83 |